Grade 6General Physics

two satellites of same mass are launched in the same orbit around the earth so as to rotate opposite to each other .if they collide inelastically and stick together as wreakage the total energy of the system just after collision is what?

ujjwal raj
12 Years agoGrade 6

1 Answer

Subhashree
7 Years ago
Here let us first consider it for one satellite...
       Potential energy U=-GMm/r
        Kinetic energy=+GMm/2r
So total energy= -GMm/r+GMm/2r
            .               = -GMm/2r
Now for two satellite ...TE=2GMm/2r
                                                   = -GMm/r.
But since it undergoes inelastic collision..mv1+mv2=(m+m)v.  
     Since they move in opposite direction v1=-v2 ..so 
Net v =0..  so no Kinetic energy only potential energy exist 
                        Thus TE=-GM(2m)/r